OpenStudy (anonymous):

Which of the following sets represents a function? {(3, 5), (-1, 7), (3, 9)} {(1, 2), (3, 2), (5, 7)} {(1, 2), (1, 4), (1, 6)}

OpenStudy (jack1):

all

OpenStudy (jack1):

what type of function specifically are you looking for?

OpenStudy (compassionate):

Wrong, Jack. It cannot b C as you can't have the same x value twice and get a different y value.

OpenStudy (compassionate):

If you look at A you see we put in 3 as the x value and got a different y value.

OpenStudy (jack1):

its a horizontal line, can be a function y = x +c

OpenStudy (anonymous):

so is it a?

OpenStudy (compassionate):

It can't be A as we use the same x-value and get different y-values It can't be C as 1 gives us differnet y values.

OpenStudy (anonymous):

oh ok i get it

OpenStudy (compassionate):

Yea, it's like if I had: y = 2 + x. If I put x = 1 y = 2 + 1. y = 3. It is impossible for me to get any other y value by inserting 1 other than 3. Basically, if the x value happens more than once and you get a different y-value, it's not a function.